# QKD security proofs for decoy-state BB84: protocol variations, proof techniques, gaps and limitations

### v1

@@ Abstract
We discuss the status of security proofs for practical decoy-state Quantum Key Distribution using the BB84 protocol, pertaining to optical implementations using weak coherent pulses and threshold photo-detectors. Our focus is on the gaps in the existing literature. Gaps might result, for example, from a mismatch of protocol detail choices and proof technique elements, from proofs relying on earlier results that made different assumptions, or from protocol choices that do not consider real-world requirements. While substantial progress has been made, our overview draws attention to the details that still require our attention.
1 Executive summary
@@
Quantum Key Distribution (QKD) is a maturing technology. From its initial conception in 1984 with the Bennett-Brassard BB84 protocol, it entered the era of experimental implementation, and soon thereafter the commercial stage as well. On the protocol side, important steps in this journey were the introduction of weak laser pulses and the introduction of the decoy state protocols. On the theory side, important steps were the development of a clean definition of the security notion of QKD protocols, and the development of proof techniques that applied initially to asymptotic security evaluations, but also to more complete proofs for finite block length of signals. Another important aspect was the consideration of the quantum optical nature of implementations that relies on optical modes rather than abstract qubits as fundamental objects of the description.
We are now in a situation where QKD appears solidly established, with well-studied protocols like the decoy-state BB84 protocol with weak coherent laser pulses. As the field is starting to define the certification environment of this application, it appears that there are many choices to be made in defining a precise QKD protocol. Moreover, full security proofs need to account for the precise choice made. Historically, proof techniques have evolved in myriad ways, with a literature that is often a patchwork of partial results. While each patch solves an important aspect of a full security proof, the patches do not necessarily fit perfectly together, either because of a mismatch in the assumptions made in various works, or because of the protocol choices that each result is applicable to.
This report starts by giving an overview of the often detailed protocol choices that are being made in different versions of the decoy-state BB84 protocol, and how those choices relate to the various proof technique elements that have been described in the literature. Based on this outline, we then give an overview of the various concepts needed for rigorous QKD security proof, before discussing common gaps and misconceptions present in the literature. Our analysis sharpens the awareness of potential pitfalls and of the level of detail that is necessary to define a QKD protocol, and to deliver a valid complete security proof. Finally, we review several important works on complete security proofs for decoy-state BB84, and list the achievements, limitations, and gaps in each of these works. In our analysis, we broadly classify security proofs into the following categories (see Section 6): entropic uncertainty relations (EUR) based proofs, phase error correction based proofs, proofs utilizing the postselection technique followed by an IID security analysis, and proofs based on the entropy accumulation theorem (EAT) or generalized entropy accumulation theorem (GEAT).
In the end, we need security proofs that
@@
  • specify the exact assumptions about the devices used in QKD,
@@
  • and deliver a mathematical proof that the given protocol with the given assumptions satisfies the exact security definition with given security parameters.
@@
  • use proof techniques that allow the extension of the protocol security to implementation security, where implementation imperfections can be taken into account.
From our overview it will become clear that currently no single publication satisfies all these criteria [1], though a tremendous progress of knowledge and insight has been achieved over the last decades. Thus, the task of assembling the knowledge to provide such a protocol definition and associated security proof is still left open.
@@ 2.1 Concepts from quantum information theory
We will also often need to refer to the notion of a quantum state conditioned on some event. We defer the formal definitions to e.g. Ref. [4, Sec. 2.2], here simply noting that this notion is formalized by requiring that the quantum state contains a classical register that determines whether the event occurred. Note that the conditional state can be defined in either a normalized or subnormalized manner — the two possibilities are often denoted as $\rho_{A|\Omega}$ and $\rho_{A\land\Omega}$ respectively (given an event $\Omega$), though this work will mainly focus on normalized conditional states, as mentioned above.
@@
The trace distance is a metric on the set of quantum states, which we shall denote as $d(\rho,\sigma)$ for any pair of quantum states $\rho,\sigma$. Here we follow the definition in e.g. Ref. [3], which covers subnormalized states. Critically, trace distance has an operational implication in terms of quantifying how well the states can be distinguished by any possible physical process, without computational limitations.
@@ 2.3 Security definitions
Current security proofs in QKD are mostly centred around a particular security definition involving trace distance. The justification behind this choice of security definition is that it yields composable security, which informally means that the keys generated from a QKD protocol can be safely used as part of a larger protocol. We defer a detailed discussion of this point to [^1] These works present somewhat different perspectives on composable security, but they coincide in terms of the desired security definitions in terms of trace distance. Refs. [5, 6]; here, we simply restate the security definitions from those works.
@@
We emphasize that the security definition does not state that the trace-distance term in Eq. 1 (for fixed-length protocols) is itself upper bounded by $\varepsilon^{\mathrm{secure}}$. This is because such an upper bound simply does not hold in general; in particular, it is usually straightforward to construct attacks such that
@@
As described in Refs. [5, 6], it is often convenient to instead study a pair of simpler conditions, referred to as correctness and secrecy: [^3] The choice to use $K_{A}$ rather than $K_{B}$ in the secrecy definition is essentially arbitrary; analogous results can be derived under either choice.
@@
[6, Theorem 2] If a QKD protocol is $\varepsilon^{\mathrm{correct}}$-correct and $\varepsilon^{\mathrm{secret}}$-secret, then it is $(\varepsilon^{\mathrm{correct}}+\varepsilon^{\mathrm{secret}})$-secure.
@@ 2.4 Completeness and expected key rate
The above definitions suffice to formulate what it means for a QKD protocol to be “secure”, in the sense that it can be “safely” used as part of a larger protocol. However, they do not fully address the question of whether the protocol is “useful”, informally speaking. For instance, observe that if a fixed-length protocol has a high output key length $\ell_{\mathrm{fixed}}$ when it accepts, but almost never actually accepts, then this would not seem to be a very practically useful protocol. [^4] While this might perhaps appear artificial, we highlight that if the conditions for the accept/abort decision are not designed with sufficient “tolerance”, then one would indeed end up with a protocol that almost always aborts. Some early works in QKD inadvertently constructed such protocols by having overly restrictive accept conditions; we briefly discuss this in Section 5.1 under the term “unique-acceptance”. More technically, inspecting Definition 2.1 shows that any fixed-length protocol in which the accept probability is never greater than $\varepsilon^{\mathrm{secure}}$ would trivially fulfill that definition, but this would not seem to be very useful either.
@@ 3 Structural elements of QKD protocols
  • An authenticated public classical channel, where Eve can see all messages sent along the channel, but cannot modify the messages without causing the authentication procedure to abort.555 Note that such a channel can be constructed from an unauthenticated public classical channel together with a short pre-shared key by using Wegman-Carter authentication [7], which has information-theoretic composable security even against quantum adversaries [8]. Hence in principle, this resource requirement could be replaced with an unauthenticated public classical channel and a short pre-shared key. However, we will not focus on this within this work. Note that this is a weaker property than the claim that Eve cannot modify the messages, and hence some care is needed when analyzing this channel in the security proof. We discuss this further in Section 5.2.
@@ 3.4.1 Error correction
This is the form of one-way error correction used in most currently implemented protocols for decoy-state BB84. Here, Alice (i.e. the party who prepared the signal states, not the one who measured them) is the party who holds the pre-amplification string after the preceding step, and the error correction procedure consists of her computing a classical string from her pre-amplification string and sending it to Bob, who then uses it together with all the data on his side to produce a guess for Alice’s pre-amplification string. Qualitatively, this is usually achieved by having Alice compute the syndrome of a suitable error-correcting code as the string to send to Bob, though there are some slight variations. For instance, some proof techniques may require Alice to furthermore apply a one-time-pad to this syndrome before sending it to Bob (see Section 6.2 for further discussion of when this is necessary), which requires a significant amount of preshared secret key to perform, though the resulting net gain in shared secret key is usually the same as other proof techniques.
@@ 3.4.2 Error verification
After error correction, the error verification procedure usually consists of one party announcing a 2-universal hash (see Section 4.2) of their pre-amplification string, with the other party checking to see whether they obtained a pre-amplification string with the same hash, and aborting if that is not the case. Qualitatively, this serves as a check on whether their pre-amplification strings match, since the properties of 2-universal hashing ensure that if the pre-amplification strings differ, the probability of their hashes matching is small. We stress however that this is a very different statement from claiming that if their hashes match, the probability of them having distinct pre-amplification strings is small. In the security proof, it would not be valid to make the latter claim; fortunately, the former claim is enough to prove that the correctness condition is satisfied, as we discuss in Section 4.3.
@@
Another slight variant of error verification is that if the subsequent privacy amplification step (Section 3.5) is also performed using 2-universal hashing, the parties can instead implement error verification after privacy amplification. To do so, one party simply announces a fixed subset of their final key bits and the other party checks if those bits match. It is straightforward to slightly modify the security proof to show that the final key with those bits excluded indeed satisfies the required correctness and secrecy properties. However, this variant may not be straightforwardly applicable if privacy amplification is performed with other extractor procedures instead.
@@ 3.5 Privacy amplification
2-universal hashing.
A possible method for privacy amplification is for one party to choose a random hash function from a 2-universal family (see Section 4.2) using some local randomness, which is sometimes termed as the seed for the privacy amplification procedure. They then publicly announce their choice, after which both parties apply that choice of hash function to their pre-amplification strings. The core theoretical tool that ensures security from this procedure is the Leftover Hashing Lemma (Section 4.4), which qualitatively states that if the input to the hash function has “enough entropy” relative to the length of the output, then the output will be nearly uniform and decoupled from any other side-information.
@@
Apart from 2-universal hashing, there are other procedures that can be used for privacy amplification, broadly referred to as randomness extractors. Qualitatively, these procedures achieve similar results to the Leftover Hashing Lemma, in that they produce outputs that are nearly uniform and decoupled from any side-information, as long as the inputs have “enough entropy” — the difference is that the operation performed on the pre-amplification string does not have to be a 2-universal hash. One noteworthy example of an extractor that does not simply apply a 2-universal hash is Trevisan’s extractor. This has the advantage that it requires less seed randomness as compared to 2-universal hashing; however, it can be more elaborate to implement, and some forms of privacy amplification theorems have only been proven for 2-universal hashing rather than Trevisan’s extractor.
@@
A technical point arises in privacy amplification regarding the domain of the input to the procedure. In particular, for some protocols the lengths of the pre-amplification strings are not fixed constants, for instance because some bits are literally discarded during the sifting procedure or other processing steps. Other protocols may instead have pre-amplification strings that are not bitstrings, for instance if bits that are sifted out are replaced with some special “blank” symbol. In such circumstances, one cannot “directly” implement procedures such as Toeplitz hashing, which were only designed to take bitstrings of fixed-length as input. A full protocol specification in such cases must explicitly state the exact procedure used for privacy amplification (for instance, a different family of Toeplitz matrices is used depending on the length of the pre-amplification strings), and the security proof must rigorously analyze this procedure accordingly. We return to this point in Section 5.6.2.

### v2

@@ Abstract
We review the current status of security proofs for practical decoy-state Quantum Key Distribution using the BB84 protocol, focusing on optical implementations with weak coherent pulses and threshold photodetectors. The primary aim of this review is to highlight gaps in the existing literature. Such gaps may arise, for instance, from mismatches between detailed protocol specifications and proof technique elements, reliance on earlier results based on different assumptions, or protocol choices that overlook real world requirements. While substantial progress has been made, our overview draws attention to the details that still demand careful attention.
1 Introduction
@@
Quantum Key Distribution (QKD) is a maturing technology. Since its initial conception in 1984 with the Bennet-Brassard BB84 [11] protocol, QKD has progressed from theoretical foundations to experimental implementations, and eventually to commercial deployment. On the protocol side, key milestones include the use of weak coherent laser pulses and the development of decoy-state methods. On the theoretically side, important advances include the formalization of a rigorous security definition for QKD protocols and the development of proof techniques – initially for asymptotic scenarios, and later extended to the scenario of finite block length of signals. A further development was the recognition of the quantum optical nature of realistic implementations that relies on optical modes rather than abstract qubits as fundamental objects of description.
1.1 Motivation
QKD is now a well-established field, with widely studied practical protocols such as the decoy-state BB84 protocol using weak coherent pulses. As the community begins to define the certification landscape of QKD technologies, it becomes clear that numerous design choices must be made in specifying a concrete QKD protocol. Importantly, complete security proofs must accurately account for these precise choices. Historically, proof techniques have evolved in myriad ways, resulting in a body of work that is often a patchwork of partial results. While each contribution addresses a critical component of a full security proof, the patches do not always integrate seamlessly – either due to mismatched assumptions across different works, or because the results apply to differing protocol specifications.
Thus, as the field moves towards certification, it is important to develop rigorous and comprehensive security proofs for practical QKD protocols. In particular, we need security proofs for practical QKD protocols that
@@ 1.1 Motivation
  • specify the exact assumptions about the devices used in the protocol,
@@
  • and deliver a mathematical proof that the given protocol with the given assumptions satisfies the given security definition with given security parameters.
@@
  • it uses proof techniques that allow the extension of the protocol security to implementation security, where implementation imperfections can be taken into account.
From this review it will become clear that currently no single publication satisfies all these criteria [1] 111 During the preparation of this manuscript, a preprint appeared on arXiv [71] presenting a promising attempt at a self-contained and rigorous security proof for a fully specified decoy-state BB84 protocol. In doing so, it attempts to meet a significantly higher standard than the works considered in this review. While this is a welcome development, this work is still undergoing scrutiny from the community, and will require time before it is thoroughly assessed. , though a tremendous progress of knowledge and insight has been achieved over the last few decades. Thus, the task of assembling the knowledge to provide such a protocol specification and associated security proof is still left open.
1.2 Focus of this review
Numerous reviews of QKD exist, covering the nature of the technology, its implementation, experimental progress, and practical devices. For example, Ref. [82] provides a pedagogical overview of the composable security framework for QKD; Ref. [124] focuses on recent advances in device-independent QKD; Ref. [122] examines imperfections in QKD implementations; and Ref. [125] offers a succinct discussion on implementation security. For foundational reviews from earlier stages of the field, see Refs. [91, 39].
The focus of this review is different. It is heavily centered on the mathematical arguments and tools required to construct a rigorous security proof for QKD. In particular, it aims to identify and analyze the existing gaps in the results pertaining to protocol security in the literature, and to highlight what would constitute a complete, modular, and verifiable proof of security for a realistic QKD protocol.
1.3 Structure of this review
This review is structured as follows: In Section 2, we set up common definitions and notations. In Section 3, we describe the QKD protocol and highlight the variety of choices that can be made in its design. These design choices directly impact how the security proof is constructed and where potential gaps may arise. In Section 4, we discuss the mathematical tools and arguments that underlie a QKD security proof, and we identify several common misconceptions and errors. Section 5 elaborates on these gaps and mistakes in existing proofs, while referring back to Section 4 for the conceptual details. In Section 6, we present a critical review of several major works on security proofs for decoy-state BB84 (Refs. [61, 88, 113, 119, 44, 49, 50, 51]). We summarize the key contributions of each work while identifying their gaps, limitations, and open issues.
There are several proof techniques currently available to theorists for establishing QKD security. These can be broadly categorized into the following four classes:
  1. 1.
    Proofs based on entropic uncertainty relations (EUR) [107],
  2. 2.
    Proofs that employ the postselection technique [75, 22] followed by a security analysis against IID collective attacks,
  3. 3.
    Phase error correction-based proofs [56],
  4. 4.
    Proofs relying on various entropy accumulation theorems (EAT) [31, 67, 32, 4].
In Section 7, we compare these proof approaches, candidly assessing their respective strengths and weaknesses (see Table 3). This comparative analysis lays the groundwork for making informed choices regarding proof techniques in future security analyses.
1.4 How to read this review
This review is written in a modular fashion and is extensively cross-referenced, allowing for multiple reading paths depending on the reader’s background and interests. Readers who are new to the field and wish to learn about security proofs are encouraged to begin with the structural elements of the QKD protocol in Section 3, and the conceptual overview in Section 4, before proceeding to the discussion of proof techniques in Section 6. Readers already familiar with security proofs may find the discussion of common gaps in Section 5, and the detailed examination of specific proofs in Section 6, the most interesting. We expect the comparison of proof technique approaches in Section 7 to be of interest to both newcomers and experts alike.
@@ 2.1 Concepts from quantum information theory
We will also often need to refer to the notion of a quantum state conditioned on some event. We defer the formal definitions to e.g. Ref. [105, Section 2.2], here simply noting that this notion is formalized by requiring that the quantum state contains a classical register that determines whether the event occurred. Note that the conditional state can be defined in either a normalized or subnormalized manner — the two possibilities are often denoted as $\rho_{A|\Omega}$ and $\rho_{A\land\Omega}$ respectively (given an event $\Omega$), though this work will mainly focus on normalized conditional states, as mentioned above.
@@
The trace distance is a metric on the set of quantum states, which we shall denote as $d(\rho,\sigma)$ for any pair of quantum states $\rho,\sigma$. Here we follow the definition in e.g. Ref. [103, Definition 3.4], which covers subnormalized states. Critically, trace distance has an operational implication in terms of quantifying how well the states can be distinguished by any possible physical process, without computational limitations.
@@ 2.3 Security definitions
Current security proofs in QKD are mostly centred around a particular security definition involving trace distance. The justification behind this choice of security definition is that it yields composable security, which informally means that the keys generated from a QKD protocol can be safely used as part of a larger protocol. [^2] A common form of such composition is if the keys generated by a QKD protocol are used for authentication in a subsequent QKD protocol. As discussed in Refs. [72, 81], a composable security framework lets us conclude that for multiple QKD protocols composed in this manner, the overall security of the entire sequence of protocols can be described in terms of the sum of the security parameters of the individual protocols. This does require the individual security parameters to be chosen sufficiently small, such that the sum of their values is still an acceptably small value. We defer a detailed discussion of this point to [^3] These works present somewhat different perspectives on composable security, but they coincide in terms of the desired security definitions in terms of trace distance. Refs. [89, 82]; here, we simply restate the security definitions from those works.
@@
We emphasize that the security definition does not state that the trace-distance term for the normalized conditional states in Eq. 1 (for fixed-length protocols) is itself upper bounded by $\varepsilon^{\mathrm{secure}}$. This is because such an upper bound simply does not hold in general; in particular, it is usually straightforward to construct attacks such that
@@
As described in Refs. [89, 82], it is often convenient to instead study a pair of simpler conditions, referred to as correctness and secrecy: [^5] The choice to use $K_{A}$ rather than $K_{B}$ in the secrecy definition is essentially a matter of convenience; analogous results can be derived under either choice.
@@
[82, Theorem 2] If a QKD protocol is $\varepsilon^{\mathrm{correct}}$-correct and $\varepsilon^{\mathrm{secret}}$-secret, then it is $(\varepsilon^{\mathrm{correct}}+\varepsilon^{\mathrm{secret}})$-secure.
@@ 2.4 Completeness and expected key rate
The above definitions suffice to formulate what it means for a QKD protocol to be “secure”, in the sense that it can be “safely” used as part of a larger protocol. However, they do not fully address the question of whether the protocol is “useful”, informally speaking. For instance, observe that if a fixed-length protocol has a high output key length $\ell_{\mathrm{fixed}}$ when it accepts, but almost never actually accepts, then this would not be a very practically useful protocol. [^6] While this might perhaps appear artificial, we highlight that if the conditions for the accept/abort decision are not designed with sufficient “tolerance”, then one would indeed end up with a protocol that almost always aborts. Some early works in QKD inadvertently constructed such protocols by having overly restrictive accept conditions; we briefly discuss this in Section 5.1 under the term “unique-acceptance”. More technically, inspecting Definition 2.1 shows that any fixed-length protocol in which the accept probability is never greater than $\varepsilon^{\mathrm{secure}}$ would trivially fulfill that definition, but this would not seem to be very useful either.
@@ 3 Structural elements of QKD protocols
In this section, we outline the various steps involved in a QKD protocol, emphasizing the range of choices available at each stage of the protocol design. These choices often have direct implications for the types of mistakes, gaps, or proof techniques discussed in later sections. Each step, along with its possible variations is explored in detail in the individual subsections.
@@
  • An authenticated public classical channel. In much of the existing QKD literature, this is assumed to mean a channel where Eve can read all messages sent along the channel, but cannot modify them. We highlight however that in practice, one can typically only construct a weaker version of this resource, in which again Eve can read all messages, but she additionally has the option of causing the recipient to abort rather than receiving the unmodified message.777 As shown in [81], it is indeed possible to obtain an information-theoretically $\varepsilon_{\text{auth}}$-secure construction (in the sense of a composable security framework) of such a resource, by using an unauthenticated public classical channel together with a short pre-shared key, via Wegman-Carter authentication [118]. Hence in principle, this resource requirement could be replaced with an unauthenticated public classical channel and a short pre-shared key. However, we will not focus on this within this work. (With this version, the natural way for the QKD protocol to handle such abort outcomes is to also abort the overall QKD protocol.) This is a weaker property than the claim that Eve cannot modify the messages, and hence some care is needed when analyzing this channel in the security proof. For now we note that the existing QKD literature, especially the usual security definition in Definition 2.1, is focused on the case where the authentication cannot abort. We discuss this further in Section 5.2.1.
@@ 3.4.1 Error correction
This is the form of one-way error correction used in most currently implemented protocols for decoy-state BB84. Here, Alice (i.e. the party who prepared the signal states, not the one who measured them) is the party who holds the pre-amplification string after the preceding step, and the error correction procedure consists of her computing a classical string from her pre-amplification string and sending it to Bob, who then uses it together with all the data on his side to produce a guess for Alice’s pre-amplification string. Qualitatively, this is usually achieved by having Alice compute the syndrome of a suitable error-correcting code as the string to send to Bob, though there are some slight variations. For instance, some proof techniques may require Alice to furthermore apply a one-time-pad to this syndrome before sending it to Bob (see Section 6.2 for further discussion of when this is necessary), which requires a significant amount of preshared secret key to perform, though the resulting net gain in shared secret key is usually the same as other proof techniques when the protocol accepts [^16] However, this approach has a notable drawback in that if the protocol aborts at some point after error correction, then the preshared key bits used to one-time-pad the error correction syndrome are already consumed and cannot be reused, resulting in a net loss of shared secret key..
@@ 3.4.2 Error verification
After error correction, the error verification procedure usually consists of one party announcing a $\delta$-AU2 hash (see Section 4.2) of their pre-amplification string, with the other party checking to see whether they obtained a pre-amplification string with the same hash, and aborting if that is not the case. Qualitatively, this serves as a check on whether their pre-amplification strings match, since the properties of $\delta$-AU2 hashing ensure that if the pre-amplification strings differ, the probability of their hashes matching is small. We stress however that this is a very different statement from claiming that if their hashes match, the probability of them having distinct pre-amplification strings is small. In the security proof, it would not be valid to make the latter claim; fortunately, the former claim is enough to prove that the correctness condition is satisfied, as we discuss in Section 4.3.
@@
Another slight variant of error verification is that if the subsequent privacy amplification step (Section 3.5) is performed using universal2 hashing (see Section 4.2), the parties can instead implement error verification after privacy amplification. To do so, one party simply announces a fixed subset of their final key bits and the other party checks if those bits match. It is straightforward to slightly modify the security proof to show that this indeed satisfies the required correctness and secrecy properties (with the announced bits being grouped into Eve’s side-information rather than the final key). However, this variant may not be straightforwardly applicable if privacy amplification is performed with other extractor procedures instead.
@@ 3.5 Privacy amplification
Universal2 hashing.
A possible method for privacy amplification is for one party to choose a random hash function according to a universal2 hashing procedure (see Section 4.2) using some local randomness, which is sometimes termed as the seed for the privacy amplification procedure. They then publicly announce their choice, after which both parties apply that choice of hash function to their pre-amplification strings. The core theoretical tool that ensures security from this procedure is the Leftover Hashing Lemma (Section 4.4), which qualitatively states that if the input to the hash function has “enough entropy” relative to the length of the output, then the output will be nearly uniform and decoupled from any other side-information.
@@
Apart from universal2 hashing, there are other procedures that can be used for privacy amplification, broadly referred to as randomness extractors. Qualitatively, these procedures achieve similar results to the Leftover Hashing Lemma, in that they produce outputs that are nearly uniform and decoupled from any side-information, as long as the inputs have “enough entropy” — the difference is that the operation performed on the pre-amplification string does not have to be a universal2 hash. One noteworthy example of an extractor that does not simply apply a universal2 hash is Trevisan’s extractor. This has the advantage that it requires less seed randomness as compared to universal2 hashing; however, it can be more elaborate to implement, and some forms of privacy amplification theorems have only been proven for universal2 hashing rather than Trevisan’s extractor.
@@
A technical point arises in privacy amplification regarding the domain of the input to the procedure. In particular, for some protocols the lengths of the pre-amplification strings are not fixed constants, for instance because some bits are literally discarded during the sifting procedure or other processing steps. Other protocols may instead have pre-amplification strings that are not bitstrings, for instance if bits that are sifted out are replaced with some special “blank” symbol. In such circumstances, one cannot “directly” implement procedures such as Toeplitz hashing, which were only designed to take bitstrings of fixed-length as input. A full protocol specification in such cases must explicitly state the exact procedure used for privacy amplification (for instance, a different family of Toeplitz matrices is used depending on the length of the pre-amplification strings), and the security proof must rigorously analyze this procedure accordingly. We return to this point in Section 5.6.2